a cyclists goes round a circular track once in 2160 seconds. If she maintains a constant speed, calculate the value of the centripetal acceleration she cycles at if the track has a radius of 15600m


1
Expert's answer
2020-08-02T15:04:37-0400

"a=v^2\/r"


"v=l/t=2\pi r/t"


"a=(2\pi r)^2/(t^2\cdot r)=4\pi^2r/t^2=4\cdot3.14^2\cdot15600/2160^2=0.132 (m/s^2)"